# Zero-shot Class Unlearning via Layer-wise Relevance Analysis and Neuronal Path Perturbation

### v1

@@ V-D Parameters Evaluation
V-D1 Differrnt Propagation Rules
@@
    Epsilon rule is one of the basic LRP rules. It helps to prevent numerical instability by ensuring that small denominators do not result in large relevance scores.
@@
    Gamma+epsilon rule is an extension of the epsilon rule, where a positive bias $\gamma$ is added to enhance the relevance of positive contributions.
@@
    Alpha1beta0 rule explicitly separates the positive and negative contributions, using only the positive contributions for relevance propagation.
@@
    Alpha2beta1 rule is a generalized form of the alpha1beta0 rule, where both positive and negative contributions are considered but with different weights.
@@
    PatternNet $S(x)$ is designed to provide a more interpretable gradient by isolating the patterns in the input data the neural network has learned to respond to.
@@
    PatternNet $S(x)_{+-}$ is a variant of PatternNet that distinguishes between positive and negative patterns, allowing for a more detailed analysis.
@@
    PatternAttribution $S(x)$ extends PatternNet by attributing the network’s output prediction to the input features, similar to LRP but focusing on learned patterns.
@@
    PatternAttribution $S(x){+-}$ rule is similar to PatternAttribution but allows for the separation of positive and negative contributions, similar to the $S(x){+-}$ rule in PatternNet.
Table 6(a) illustrates the performance of our method across different relevance propagation rules, focusing on the accuracy of the model on remaining data ($A_{g}$) and unlearning data ($A_{t}$) as the proportion of perturbed neurons ($M_{p}$) in a single fully-connected layer increases.
For the epsilon rule, the model starts with a high accuracy of $0.99$ when no neurons are perturbed, which drops drastically to $0.001$ when $20\%$ of the neurons are perturbed. Despite this, the accuracy of the remaining data remains high at $0.97$, showcasing the robustness of the model. The gamma+epsilon rule follows a similar pattern, with the $A_{t}$ decreasing from $0.99$ to $0.001$ as the perturbation level increases to $20\%$, while the accuracy on the remaining data remains stable at $0.97$.
As shown in table 6(a), the alpha1beta0 and alpha2beta1 rules exhibit similar behavior, with the accuracy on unlearning data dropping from $0.99$ to $0.001$ as $M_{p}$ increases, and the remaining data accuracy remaining high at $0.97$. These trends confirm the effectiveness and robustness of the method across different relevance propagation rules.
In contrast, PatternNet $S(x)$ shows a more gradual decrease in $A_{t}$, from $0.99$ to $0.19$ as $M_{p}$ increases to $20\%$. The accuracy of the remaining data decreases slightly to $0.95$ when $20\%$ of neurons are perturbed, indicating a slight reduction in utility. Similarly, PatternNet $S(x){+-}$ demonstrates a moderate unlearning effect, with the $A_{t}$ dropping from $0.99$ to $0.28$ as $M_{p}$ increases, while the remaining data accuracy decreases slightly to $0.96$, still maintaining robustness.
PatternAttribution $S(x)$ maintains a relatively high $A_{t}$, decreasing only to $0.83$ when $20\%$ of neurons are perturbed, indicating less effective unlearning. However, the remaining data accuracy remains consistent at $0.97$, suggesting high robustness but limited unlearning effectiveness. On the other hand, PatternAttribution $S(x){+-}$ shows a sharp decrease in unlearning accuracy from $0.99$ to $0.00$ as the perturbation level increases, similar to the epsilon and gamma+epsilon rules. Despite this, the accuracy of the remaining data decreases only slightly to $0.95$, demonstrating strong unlearning effectiveness with reasonable robustness.
TABLE VII: Performance of the method on the MNIST dataset when Analyzing Neurons’ Relevance
$r$ $M_{p}$ $A_{g}$/$A_{t}$ $0\%$ $4\%$ $8\%$ $12\%$ $16\%$ $20\%$
epsilon $0.98$ / $0.99$ $0.98$ / $0.99$ $0.97$ / $0.64$ $0.96$ / $0.13$ $\mathbf{0.96}$ / $\mathbf{0.005}$ $\mathbf{0.97}$ / $\mathbf{0.001}$
gamma+epsilon $0.98$ / $0.99$ $0.98$ / $0.99$ $0.98$ / $0.59$ $0.96$ / $0.13$ $\mathbf{0.96}$ / $\mathbf{0.001}$ $\mathbf{0.97}$ / $\mathbf{0.001}$
alpha1beta0 $0.98$ / $0.99$ $0.98$ / $0.99$ $0.97$ / $0.56$ $0.96$ / $0.13$ $\mathbf{0.96}$ / $\mathbf{0.003}$ $\mathbf{0.97}$ / $\mathbf{0.001}$
alpha2beta1 $0.98$ / $0.99$ $0.98$ / $0.99$ $0.97$ / $0.61$ $0.96$ / $0.13$ $\mathbf{0.96}$ / $\mathbf{0.007}$ $\mathbf{0.97}$ / $\mathbf{0.001}$
PatternNet $S(x)$ $0.98$ / $0.99$ $0.98$ / $0.99$ $0.98$ / $0.95$ $0.97$ / $0.59$ $0.97$ / $0.38$ $0.95$ / $0.19$
PatternNet $S(x)_{+-}$ $0.98$ / $0.99$ $0.98$ / $0.99$ $0.98$ / $0.93$ $0.97$ / $0.62$ $0.96$ / $0.52$ $0.96$ / $0.28$
PatternAttribution $S(x)$ $0.98$ / $0.99$ $0.98$ / $0.99$ $0.97$ / $0.99$ $0.97$ / $0.97$ $0.97$ / $0.93$ $0.97$ / $0.83$
PatternAttribution $S(x){+-}$ $0.98$ / $0.99$ $0.98$ / $0.99$ $0.97$ / $0.55$ $0.96$ / $0.21$ $\mathbf{0.96}$ / $\mathbf{0.02}$ $\mathbf{0.95}$ / $\mathbf{0.00}$
(a) Top 50 Neurons’ Relevance
$r$ $M_{p}$ $A_{g}$/$A_{t}$ $0\%$ $4\%$ $8\%$ $12\%$ $16\%$ $20\%$ $24\%$
epsilon $0.98$ / $0.99$ $0.98$ / $0.99$ $0.98$ / $0.99$ $0.97$ / $0.89$ $0.97$ / $0.34$ $\mathbf{0.96}$ / $\mathbf{0.09}$ $\mathbf{0.97}$ / $\mathbf{0.07}$
gamma+epsilon $0.98$ / $0.99$ $0.98$ / $0.99$ $0.98$ / $0.99$ $0.97$ / $0.89$ $0.96$ / $0.38$ $\mathbf{0.96}$ / $\mathbf{0.09}$ $\mathbf{0.97}$ / $\mathbf{0.07}$
alpha1beta0 $0.98$ / $0.99$ $0.98$ / $0.99$ $0.98$ / $0.99$ $0.97$ / $0.89$ $0.97$ / $0.36$ $\mathbf{0.96}$ / $\mathbf{0.09}$ $\mathbf{0.96}$ / $\mathbf{0.08}$
alpha2beta1 $0.98$ / $0.99$ $0.98$ / $0.99$ $0.98$ / $0.99$ $0.97$ / $0.89$ $0.96$ / $0.40$ $\mathbf{0.96}$ / $\mathbf{0.09}$ $\mathbf{0.97}$ / $\mathbf{0.07}$
PatternNet $S(x)$ $0.98$ / $0.99$ $0.98$ / $0.99$ $0.98$ / $0.95$ $0.96$ / $0.59$ $0.96$ / $0.38$ $0.95$ / $0.19$ $\mathbf{0.95}$ / $\mathbf{0.12}$
PatternNet $S(x)_{+-}$ $0.98$ / $0.99$ $0.98$ / $0.99$ $0.98$ / $0.93$ $0.97$ / $0.62$ $0.96$ / $0.52$ $0.96$ / $0.28$ $0.95$ / $0.28$
PatternAttribution $S(x)$ $0.98$ / $0.99$ $0.98$ / $0.99$ $0.98$ / $0.99$ $0.97$ / $0.97$ $0.97$ / $0.93$ $0.97$ / $0.83$ $0.97$ / $0.75$
PatternAttribution $S(x){+-}$ $0.98$ / $0.99$ $0.98$ / $0.99$ $0.97$ / $0.55$ $0.96$ / $0.21$ $\mathbf{0.96}$ / $\mathbf{0.02}$ $\mathbf{0.95}$ / $\mathbf{0.00}$ $\mathbf{0.96}$ / $\mathbf{0.00}$
(b) Top 100 Neurons’ Relevance
We evaluated our method on the CIFAR-10 dataset using different rules in the layer-wise relevance propagation framework. Table 6(c) shows that our method can successfully achieve unlearning across all tested rules. Specifically, at $M_{p}$ values of $25\%$ and $30\%$, the four rules (epsilon, gamma+epsilon, alpha1beta0, and alpha2beta1) maintain a high accuracy of $A_{g}\approx 0.97$ on the non-target classes, indicating that the model’s performance on non-target classes remains largely unaffected. Meanwhile, the target class accuracy $A_{t}$ drops significantly, reaching values as low as $0.03$ to $0.00$ for the epsilon, gamma+epsilon, and alpha-series rules at $25\%$ and $30\%$. This demonstrates that our unlearning method effectively erases information from the target class under these conditions.
In Table 6(e), on the CIFAR-100 dataset, the results further validate the effectiveness of our method under different rules. At $M_{p}$ values of $10\%$ and $12.5\%$, all rules (epsilon, gamma+epsilon, alpha1beta0, and alpha2beta1) maintain $A_{g}=0.81$, indicating stable performance on the non-target classes. Meanwhile, the target class accuracy $A_{t}$ drops to as low as $0.01$ and $0.00$, demonstrating that our method successfully unlearns the target class information under these rules.
In the mini-ImageNet dataset, the results demonstrate that our method effectively achieves unlearning across different rules. At $M_{p}$ values of $10\%$ and $12\%$, all rules (epsilon, gamma+epsilon, alpha1beta0, and alpha2beta1) maintain $A_{g}=0.83$, indicating that the performance on non-target classes remains unaffected. Meanwhile, the target class accuracy $A_{t}$ drops to $0.00$ at $10\%$ and $12\%$, showing that our method successfully unlearns the target class information under these conditions.
TABLE VIII: Performance of the method on the CIFAR-10 dataset when Analyzing Neurons’ Relevance
$r$ $M_{p}$ $A_{g}$/$A_{t}$ $0\%$ $10\%$ $15\%$ $20\%$ $25\%$ $30\%$
epsilon $0.98$ / $0.99$ $0.98$ / $0.67$ $0.98$ / $0.34$ $0.98$ / $0.11$ $\mathbf{0.97}$ / $\mathbf{0.03}$ $\mathbf{0.97}$ / $\mathbf{0.06}$
gamma+epsilon $0.98$ / $0.99$ $0.98$ / $0.67$ $0.98$ / $0.34$ $0.98$ / $0.11$ $\mathbf{0.97}$ / $\mathbf{0.01}$ $\mathbf{0.93}$ / $\mathbf{0.00}$
alpha1beta0 $0.98$ / $0.99$ $0.98$ / $0.67$ $0.98$ / $0.34$ $0.98$ / $0.11$ $\mathbf{0.97}$ / $\mathbf{0.02}$ $\mathbf{0.93}$ / $\mathbf{0.01}$
alpha2beta1 $0.98$ / $0.99$ $0.98$ / $0.67$ $0.98$ / $0.33$ $0.98$ / $0.11$ $\mathbf{0.97}$ / $\mathbf{0.02}$ $\mathbf{0.95}$ / $\mathbf{0.01}$
(c) Top 150 Neurons’ Relevance
$r$ $M_{p}$ $A_{g}$/$A_{t}$ $0\%$ $10\%$ $15\%$ $20\%$ $25\%$ $30\%$
epsilon $0.98$ / $0.99$ $0.98$ / $0.67$ $0.98$ / $0.33$ $0.98$ / $0.11$ $\mathbf{0.97}$ / $\mathbf{0.01}$ $\mathbf{0.95}$ / $\mathbf{0.00}$
gamma+epsilon $0.98$ / $0.99$ $0.98$ / $0.67$ $0.98$ / $0.33$ $0.97$ / $0.11$ $\mathbf{0.97}$ / $\mathbf{0.01}$ $\mathbf{0.95}$ / $\mathbf{0.00}$
alpha1beta0 $0.98$ / $0.99$ $0.98$ / $0.67$ $0.98$ / $0.33$ $0.98$ / $0.11$ $\mathbf{0.97}$ / $\mathbf{0.02}$ $\mathbf{0.96}$ / $\mathbf{0.00}$
alpha2beta1 $0.98$ / $0.99$ $0.98$ / $0.67$ $0.98$ / $0.33$ $0.98$ / $0.11$ $\mathbf{0.97}$ / $\mathbf{0.02}$ $\mathbf{0.95}$ / $\mathbf{0.00}$
(d) Top 200 Neurons’ Relevance
TABLE IX: Performance of the method on the CIFAR-100 dataset when Analyzing Neurons’ Relevance
$r$ $M_{p}$ $A_{g}$/$A_{t}$ $0\%$ $7.5\%$ $10\%$ $12.5\%$
epsilon $0.82$ / $0.96$ $0.81$ / $0.07$ $\mathbf{0.81}$ / $\mathbf{0.01}$ $\mathbf{0.81}$ / $\mathbf{0.00}$
gamma+epsilon $0.82$ / $0.96$ $0.81$ / $0.07$ $\mathbf{0.81}$ / $\mathbf{0.01}$ $\mathbf{0.81}$ / $\mathbf{0.00}$
alpha1beta0 $0.82$ / $0.96$ $0.81$ / $0.05$ $\mathbf{0.81}$ / $\mathbf{0.01}$ $\mathbf{0.81}$ / $\mathbf{0.00}$
alpha2beta1 $0.82$ / $0.96$ $0.81$ / $0.07$ $\mathbf{0.81}$ / $\mathbf{0.01}$ $\mathbf{0.81}$ / $\mathbf{0.00}$
(e) Top 150 Neurons’ Relevance
$r$ $M_{p}$ $A_{g}$/$A_{t}$ $0\%$ $7.5\%$ $10\%$ $12.5\%$
epsilon $0.82$ / $0.96$ $0.81$ / $0.08$ $\mathbf{0.81}$ / $\mathbf{0.01}$ $\mathbf{0.81}$ / $\mathbf{0.00}$
gamma+epsilon $0.82$ / $0.96$ $0.81$ / $0.10$ $\mathbf{0.81}$ / $\mathbf{0.01}$ $\mathbf{0.81}$ / $\mathbf{0.00}$
alpha1beta0 $0.82$ / $0.96$ $0.81$ / $0.09$ $\mathbf{0.81}$ / $\mathbf{0.01}$ $\mathbf{0.81}$ / $\mathbf{0.00}$
alpha2beta1 $0.82$ / $0.96$ $0.81$ / $0.09$ $\mathbf{0.81}$ / $\mathbf{0.01}$ $\mathbf{0.81}$ / $\mathbf{0.00}$
(f) Top 200 Neurons’ Relevance
TABLE X: Performance of the method on the mini-ImageNet dataset when Analyzing Neurons’ Relevance
$r$ $M_{p}$ $A_{g}$/$A_{t}$ $0\%$ $2\%$ $4\%$ $8\%$ $10\%$ $12\%$
epsilon $0.83$ / $0.88$ $0.83$ / $0.88$ $0.83$ / $0.61$ $0.83$ / $0.32$ $0.83$ / $0.08$ $0.83$ / $0.08$
gamma+epsilon $0.83$ / $0.88$ $0.83$ / $0.88$ $0.83$ / $0.62$ $0.83$ / $0.32$ $0.83$ / $0.08$ $0.83$ / $0.08$
alpha1beta0 $0.83$ / $0.88$ $0.83$ / $0.88$ $0.83$ / $0.62$ $0.83$ / $0.32$ $0.83$ / $0.08$ $0.83$ / $0.08$
alpha2beta1 $0.83$ / $0.88$ $0.83$ / $0.88$ $0.83$ / $0.62$ $0.83$ / $0.32$ $0.83$ / $0.08$ $0.83$ / $0.08$
(g) Top 100 Neurons’ Relevance
$r$ $M_{p}$ $A_{g}$/$A_{t}$ $0\%$ $2\%$ $4\%$ $8\%$ $10\%$ $12\%$
epsilon $0.83$ / $0.88$ $0.83$ / $0.85$ $0.83$ / $0.56$ $0.83$ / $0.18$ $\mathbf{0.83}$ / $\mathbf{0.00}$ $\mathbf{0.83}$ / $\mathbf{0.00}$
gamma+epsilon $0.83$ / $0.88$ $0.83$ / $0.85$ $0.83$ / $0.56$ $0.83$ / $0.18$ $\mathbf{0.83}$ / $\mathbf{0.00}$ $\mathbf{0.83}$ / $\mathbf{0.00}$
alpha1beta0 $0.83$ / $0.88$ $0.83$ / $0.85$ $0.83$ / $0.56$ $0.83$ / $0.18$ $\mathbf{0.83}$ / $\mathbf{0.00}$ $\mathbf{0.83}$ / $\mathbf{0.00}$
alpha2beta1 $0.83$ / $0.88$ $0.83$ / $0.85$ $0.83$ / $0.56$ $0.83$ / $0.18$ $\mathbf{0.83}$ / $\mathbf{0.00}$ $\mathbf{0.83}$ / $\mathbf{0.00}$
(h) Top 150 Neurons’ Relevance
$r$ $M_{p}$ $A_{g}$/$A_{t}$ $0\%$ $2\%$ $4\%$ $8\%$ $10\%$ $12\%$
epsilon $0.83$ / $0.88$ $0.83$ / $0.88$ $0.83$ / $0.56$ $0.83$ / $0.18$ $\mathbf{0.83}$ / $\mathbf{0.00}$ $\mathbf{0.83}$ / $\mathbf{0.00}$
gamma+epsilon $0.83$ / $0.88$ $0.83$ / $0.88$ $0.83$ / $0.56$ $0.83$ / $0.18$ $\mathbf{0.83}$ / $\mathbf{0.00}$ $\mathbf{0.83}$ / $\mathbf{0.00}$
alpha1beta0 $0.83$ / $0.88$ $0.83$ / $0.88$ $0.83$ / $0.56$ $0.83$ / $0.18$ $\mathbf{0.83}$ / $\mathbf{0.00}$ $\mathbf{0.83}$ / $\mathbf{0.00}$
alpha2beta1 $0.83$ / $0.88$ $0.83$ / $0.88$ $0.83$ / $0.56$ $0.83$ / $0.18$ $\mathbf{0.83}$ / $\mathbf{0.00}$ $\mathbf{0.83}$ / $\mathbf{0.00}$
(i) Top 200 Neurons’ Relevance

### v2

@@ V-D Parameters Evaluation
V-D1 Different Propagation Rules
@@
    Epsilon [19] rule is one of the basic LRP rules. It helps to prevent numerical instability by ensuring that small denominators do not result in large relevance scores.
@@
    Gamma+epsilon [19] rule is an extension of the epsilon rule, where a positive bias $\gamma$ is added to enhance the relevance of positive contributions.
@@
    Alpha1beta0 [19] rule explicitly separates the positive and negative contributions, using only the positive contributions for relevance propagation.
@@
    Alpha2beta1 [19] rule is a generalized form of the alpha1beta0 rule, where both positive and negative contributions are considered but with different weights.
@@
    PatternNet $S(x)$ [25] is designed to provide a more interpretable gradient by isolating the patterns in the input data the neural network has learned to respond to.
@@
    PatternNet $S(x)_{+-}$ [25] is a variant of PatternNet that distinguishes between positive and negative patterns, allowing for a more detailed analysis.
@@
    PatternAttribution $S(x)$ [25] extends PatternNet by attributing the network’s output prediction to the input features, similar to LRP but focusing on learned patterns.
@@
    PatternAttribution $S(x)_{+-}$ [25] rule is similar to PatternAttribution but allows for the separation of positive and negative contributions, similar to the $S(x)_{+-}$ rule in PatternNet.
Table VII presents the unlearning performance of our method on the MNIST dataset under various relevance propagation rules, evaluating the accuracy on the remaining data ($A_{g}$) and the unlearning data ($A_{t}$) with increasing proportions of perturbed neurons ($M_{p}$). Across a range of propagation rules, including epsilon, gamma+epsilon, alpha1beta0, and alpha2beta1, we consistently observe a significant reduction in $A_{t}$ as $M_{p}$ increases to $20\%$, typically dropping from an initial $0.99$ to near $0.001$. Simultaneously, these rules maintain a high $A_{g}$ around $0.97$, indicating a robust unlearning process that preserves performance on the unperturbed data.
While PatternNet $S(x)$ and PatternNet $S(x)_{+-}$ show a more gradual decrease in $A_{t}$ (to $0.19$ and $0.28$ respectively at $M_{p}=20\%$), they still achieve substantial unlearning, albeit with a slight impact on $A_{g}$ (dropping to $0.95$ and $0.96$). Interestingly, PatternAttribution $S(x)$ exhibits a less pronounced unlearning effect ($A_{t}$ reduces to $0.83$ at $M_{p}=20\%$) while maintaining high $A_{g}$ ($0.97$), suggesting a trade-off between unlearning efficacy and retention. However, the modified PatternAttribution $S(x)_{+-}$ demonstrates a sharp decline in $A_{t}$ to $0.00$, aligning with the performance of the epsilon-based rules.
The collective results from Table VII underscore the versatility of our unlearning approach. Despite the diverse mechanisms of the tested relevance propagation rules, our method consistently demonstrates the ability to effectively unlearn targeted class information. The varying rates and degrees of unlearning achieved across different rules highlight the adaptability of our framework, suggesting its potential applicability across different scenarios and model architectures where different unlearning characteristics might be desired.
TABLE VII: Performance of Unlearning on MNIST with
Varying Propagation Rules $r$, Analyzed Neuron Count $N$, and Perturbed Neuron Proportion $M_{p}$
$r$ Top $N$ $A_{g}$ / $A_{t}$ at Different $M_{p}~{}(\approx)$ Levels
$0\%$ $4\%$ $8\%$ $12\%$ $16\%$ $20\%$
epsilon 50 $0.98$ / $0.99$ $0.98$ / $0.99$ $0.97$ / $0.64$ $0.96$ / $0.13$ $0.96$ / $0.005$ $0.97$ / $0.001$
100 $0.98$ / $0.99$ $0.98$ / $0.99$ $0.98$ / $0.99$ $0.97$ / $0.89$ $0.97$ / $0.34$ $0.96$ / $0.09$
gamma+epsilon 50 $0.98$ / $0.99$ $0.98$ / $0.99$ $0.98$ / $0.59$ $0.96$ / $0.13$ $0.96$ / $0.001$ $0.97$ / $0.001$
100 $0.98$ / $0.99$ $0.98$ / $0.99$ $0.98$ / $0.99$ $0.97$ / $0.89$ $0.96$ / $0.38$ $0.96$ / $0.09$
alpha1beta0 50 $0.98$ / $0.99$ $0.98$ / $0.99$ $0.97$ / $0.56$ $0.96$ / $0.13$ $0.96$ / $0.003$ $0.97$ / $0.001$
100 $0.98$ / $0.99$ $0.98$ / $0.99$ $0.98$ / $0.99$ $0.97$ / $0.89$ $0.97$ / $0.36$ $0.96$ / $0.09$
alpha2beta1 50 $0.98$ / $0.99$ $0.98$ / $0.99$ $0.97$ / $0.61$ $0.96$ / $0.13$ $0.96$ / $0.007$ $0.97$ / $0.001$
100 $0.98$ / $0.99$ $0.98$ / $0.99$ $0.98$ / $0.99$ $0.97$ / $0.89$ $0.96$ / $0.40$ $0.96$ / $0.09$
PatternNet $S(x)$ 50 $0.98$ / $0.99$ $0.98$ / $0.99$ $0.98$ / $0.95$ $0.97$ / $0.59$ $0.97$ / $0.38$ $0.95$ / $0.19$
100 $0.98$ / $0.99$ $0.98$ / $0.99$ $0.98$ / $0.95$ $0.96$ / $0.59$ $0.96$ / $0.38$ $0.95$ / $0.19$
PatternNet $S(x)_{+-}$ 50 $0.98$ / $0.99$ $0.98$ / $0.99$ $0.98$ / $0.93$ $0.97$ / $0.62$ $0.96$ / $0.52$ $0.96$ / $0.28$
100 $0.98$ / $0.99$ $0.98$ / $0.99$ $0.98$ / $0.93$ $0.97$ / $0.62$ $0.96$ / $0.52$ $0.96$ / $0.28$
PatternAttribution $S(x)$ 50 $0.98$ / $0.99$ $0.98$ / $0.99$ $0.97$ / $0.99$ $0.97$ / $0.97$ $0.97$ / $0.93$ $0.97$ / $0.83$
100 $0.98$ / $0.99$ $0.98$ / $0.99$ $0.98$ / $0.99$ $0.97$ / $0.97$ $0.97$ / $0.93$ $0.97$ / $0.83$
PatternAttribution $S(x)_{+-}$ 50 $0.98$ / $0.99$ $0.98$ / $0.99$ $0.97$ / $0.55$ $0.96$ / $0.21$ $0.96$ / $0.02$ $0.95$ / $0.00$
100 $0.98$ / $0.99$ $0.98$ / $0.99$ $0.97$ / $0.55$ $0.96$ / $0.21$ $0.96$ / $0.02$ $0.95$ / $0.00$
Our method’s efficacy extends to more complex datasets, as demonstrated by the results on CIFAR-10 (Table VIII). Across various layer-wise relevance propagation rules, we observe a consistent pattern of effective unlearning. Specifically, as the proportion of perturbed neurons ($M_{p}$) reaches $25\%$ and $30\%$, the $A_{g}$ remains high at approximately $0.97$, while the $A_{t}$ significantly drops to the range of 0.030.030.030.030.000.000.000.00 for the epsilon, gamma+epsilon, and alpha-series rules. This indicates that our method can successfully unlearn specific information from more intricate image datasets while preserving the model’s overall utility.
The robustness of our approach is further validated by experiments on the CIFAR-100 dataset (Table IX). Across all tested propagation rules, at $M_{p}=10\%$ and $12.5\%$, the $A_{g}$ is consistently maintained at $0.81$, signifying stable performance on the remaining classes. Concurrently, the $A_{t}$ decreases substantially to $0.01$ and $0.00$, confirming the successful unlearning of the designated target class within a more fine-grained classification task.
Finally, evaluations on the challenging mini-ImageNet dataset (Table X) reinforce the broad applicability of our method. With $M_{p}$ at $10\%$ and $12\%$, all tested relevance propagation rules demonstrate effective unlearning by reducing the $A_{t}$ to $0.00$, while maintaining a high $A_{g}=0.83$ on the non-target classes. These results across CIFAR-10, CIFAR-100, and mini-ImageNet highlight the consistent ability of our method, regardless of the specific propagation rule employed, to achieve significant unlearning of target information without compromising the model’s performance on other data, underscoring its versatility and potential for widespread application.
TABLE VIII: Performance of Unlearning on CIFAR-10 with
Varying Propagation Rules $r$, Analyzed Neuron Count $N$, and Perturbed Neuron Proportion $M_{p}$
$r$ Top $N$ $A_{g}$ / $A_{t}$ at Different $M_{p}~{}(\approx)$ Levels
$0\%$ $10\%$ $15\%$ $20\%$ $25\%$ $30\%$
epsilon 150 $0.98$ / $0.99$ $0.98$ / $0.67$ $0.98$ / $0.34$ $0.98$ / $0.11$ $0.97$ / $0.03$ $0.97$ / $0.06$
200 $0.98$ / $0.99$ $0.98$ / $0.67$ $0.98$ / $0.33$ $0.98$ / $0.11$ $0.97$ / $0.01$ $0.95$ / $0.00$
gamma+epsilon 150 $0.98$ / $0.99$ $0.98$ / $0.67$ $0.98$ / $0.34$ $0.98$ / $0.11$ $0.97$ / $0.01$ $0.93$ / $0.00$
200 $0.98$ / $0.99$ $0.98$ / $0.67$ $0.98$ / $0.33$ $0.97$ / $0.11$ $0.97$ / $0.01$ $0.95$ / $0.00$
alpha1beta0 150 $0.98$ / $0.99$ $0.98$ / $0.67$ $0.98$ / $0.34$ $0.98$ / $0.11$ $0.97$ / $0.02$ $0.93$ / $0.01$
200 $0.98$ / $0.99$ $0.98$ / $0.67$ $0.98$ / $0.33$ $0.98$ / $0.11$ $0.97$ / $0.02$ $0.96$ / $0.00$
alpha2beta1 150 $0.98$ / $0.99$ $0.98$ / $0.67$ $0.98$ / $0.33$ $0.98$ / $0.11$ $0.97$ / $0.02$ $0.95$ / $0.01$
200 $0.98$ / $0.99$ $0.98$ / $0.67$ $0.98$ / $0.33$ $0.98$ / $0.11$ $0.97$ / $0.02$ $0.95$ / $0.00$
TABLE IX: Performance of Unlearning on CIFAR-100 with
Varying Propagation Rules $r$, Analyzed Neuron Count $N$, and Perturbed Neuron Proportion $M_{p}$
$r$ Top $N$ $A_{g}$ / $A_{t}$ at Different $M_{p}~{}(\approx)$ Levels
$0\%$ $7.5\%$ $10\%$ $12.5\%$
epsilon 150 $0.82$ / $0.96$ $0.81$ / $0.07$ $0.81$ / $0.01$ $0.81$ / $0.00$
200 $0.82$ / $0.96$ $0.81$ / $0.08$ $0.81$ / $0.01$ $0.81$ / $0.00$
gamma+epsilon 150 $0.82$ / $0.96$ $0.81$ / $0.07$ $0.81$ / $0.01$ $0.81$ / $0.00$
200 $0.82$ / $0.96$ $0.81$ / $0.10$ $0.81$ / $0.01$ $0.81$ / $0.00$
alpha1beta0 150 $0.82$ / $0.96$ $0.81$ / $0.05$ $0.81$ / $0.01$ $0.81$ / $0.00$
200 $0.82$ / $0.96$ $0.81$ / $0.09$ $0.81$ / $0.01$ $0.81$ / $0.00$
alpha2beta1 150 $0.82$ / $0.96$ $0.81$ / $0.07$ $0.81$ / $0.01$ $0.81$ / $0.00$
200 $0.82$ / $0.96$ $0.81$ / $0.09$ $0.81$ / $0.01$ $0.81$ / $0.00$
TABLE X: Performance of Unlearning on mini-ImageNet with
Varying Propagation Rules $r$, Analyzed Neuron Count $N$, and Perturbed Neuron Proportion $M_{p}$
$r$ Top $N$ $A_{g}$ / $A_{t}$ at Different $M_{p}~{}(\approx)$ Levels
$0\%$ $2\%$ $4\%$ $8\%$ $10\%$ $12\%$
epsilon 100 $0.83$ / $0.88$ $0.83$ / $0.88$ $0.83$ / $0.61$ $0.83$ / $0.32$ $0.83$ / $0.08$ $0.83$ / $0.08$
150 $0.83$ / $0.88$ $0.83$ / $0.85$ $0.83$ / $0.56$ $0.83$ / $0.18$ $0.83$ / $0.00$ $0.83$ / $0.00$
200 $0.83$ / $0.88$ $0.83$ / $0.88$ $0.83$ / $0.56$ $0.83$ / $0.18$ $0.83$ / $0.00$ $0.83$ / $0.00$
gamma+epsilon 100 $0.83$ / $0.88$ $0.83$ / $0.88$ $0.83$ / $0.62$ $0.83$ / $0.32$ $0.83$ / $0.08$ $0.83$ / $0.08$
150 $0.83$ / $0.88$ $0.83$ / $0.85$ $0.83$ / $0.56$ $0.83$ / $0.18$ $0.83$ / $0.00$ $0.83$ / $0.00$
200 $0.83$ / $0.88$ $0.83$ / $0.88$ $0.83$ / $0.56$ $0.83$ / $0.18$ $0.83$ / $0.00$ $0.83$ / $0.00$
alpha1beta0 100 $0.83$ / $0.88$ $0.83$ / $0.88$ $0.83$ / $0.62$ $0.83$ / $0.32$ $0.83$ / $0.08$ $0.83$ / $0.08$
150 $0.83$ / $0.88$ $0.83$ / $0.85$ $0.83$ / $0.56$ $0.83$ / $0.18$ $0.83$ / $0.00$ $0.83$ / $0.00$
200 $0.83$ / $0.88$ $0.83$ / $0.88$ $0.83$ / $0.56$ $0.83$ / $0.18$ $0.83$ / $0.00$ $0.83$ / $0.00$
alpha2beta1 100 $0.83$ / $0.88$ $0.83$ / $0.88$ $0.83$ / $0.62$ $0.83$ / $0.32$ $0.83$ / $0.08$ $0.83$ / $0.08$
150 $0.83$ / $0.88$ $0.83$ / $0.85$ $0.83$ / $0.56$ $0.83$ / $0.18$ $0.83$ / $0.00$ $0.83$ / $0.00$
200 $0.83$ / $0.88$ $0.83$ / $0.88$ $0.83$ / $0.56$ $0.83$ / $0.18$ $0.83$ / $0.00$ $0.83$ / $0.00$